Magnesium has three stable isotopes: 24 Mg, 25 Mg, and 26 Mg. The lightest isotope has an abundance of 78.90 %. Calculate the percentage abundance of the other isotopes.


1
Expert's answer
2021-11-02T02:06:59-0400

Isotopic mass of 24Mg= 23.9850amu

Isotopic mass of 25Mg=24.9858amu

Isotopic mass of 26Mg=25.9826amu

Atomic mass of Mg=24.3050amu


Isotopic abundance of 24Mg=78.90%

Isotopic abundance of 25Mg and 26Mg;


=100%-78.90%

=21.1%


Let isotopic abundance of 25Mg= x%

Therefore isotopic abundance of 26Mg

=(21.1 -x)%


="\frac {78.90}{100} (23.9850)+\frac{x}{100}(24.9858)"


"+\frac{21.1-x}{100}(25.9826)=24.3050"


Solving for x;


"0.9968x=10.1494"

"x=10.18%" "\%"

"21.1-x=21.1-10.198"

"10.92\%"



Isotopic abundance of "25Mg=10.18\%"


Isotopic abundance of"\>26Mg=10.92\%"